OpenStudy (anonymous):

Perform the indicated operations and simplify. [x + (7 + x2)][x − (7 + x2)]

OpenStudy (anonymous):

Use (A - B)(A + B) = A^2 - B^2

OpenStudy (anonymous):

Here A = x and B = (......)

OpenStudy (anonymous):

(7+x^2)

OpenStudy (anonymous):

\[x^2-49+14x^2+x^4\] .... is this correct?

OpenStudy (anonymous):

wait it should be 15x^2 right?

OpenStudy (anonymous):

Almost but no

OpenStudy (anonymous):

15 is dead wrong

OpenStudy (anonymous):

Simply open the brackets with minus sign more methoducally - u need more minuses

OpenStudy (anonymous):

\[x^2-49-14x^2-x^4\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Buuuut it is NOT the end (nothing is, but that's mystics) of computation as required by the question

OpenStudy (anonymous):

Thanks for the help Mikael. I have some other concepts I'm still brushing up on if you're willing to help correct those as well. I understand if you can't; either way thanks for helping out.

OpenStudy (anonymous):

Yea I will gladly help but same time TOMORROW (it can be 20 minutes just for ur questions)

OpenStudy (anonymous):

Anyway - here GATHER the SIMILAR TERMS - this is called "simplify"

OpenStudy (anonymous):

Aaaand don't forget to "fan" me

OpenStudy (anonymous):

So you were saying this can be simplified even further?

OpenStudy (anonymous):

Yep that's correct

OpenStudy (anonymous):

alright I think I found it through factoring

OpenStudy (anonymous):

Gather Like terms (this should be said in a "voice from heavens" reverberating echo low bass pitch :))) )

OpenStudy (anonymous):

\[(x-7)(x+7)-x^2(14-x^2)\]

OpenStudy (anonymous):

No No No - this is ANTI-SIMPLIFYING !!!

OpenStudy (anonymous):

It is smart and inventive but they want a veryy dumm GATHER the LIKE-TERMS ...

OpenStudy (anonymous):

-49-15x^2-4x

OpenStudy (anonymous):

\[x^2 - 14x^2 = -13x^2\]

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

\[-x^4 -13x^2 - 49\]

OpenStudy (anonymous):

This IS the answer. God Bless and good Bye